OpenStudy (anonymous):

Samantha is going on vacation for the summer and is trying to choose between two different plans. The first plan costs $450 for 3 days at a hotel and 2 days at an amusement park. The second plan offers 5 days at the same hotel and 3 days at the amusement park for $700. The cost of 1 day at the hotel and the amusement park is the same under both plans. How much does a 1 day trip to the amusement park cost?

OpenStudy (anonymous):

Define h=cost of hotel per day a=cost of amusement park per day Then set-up and solve the system: 3h+2a=450 [1] 5h+3a=700 [2] Mutliply [1] by 3 and mult [2] by -2, and make a linear combination to eliminate the a-variables: 9h+6a=1350 [1]' -10h-6a=-1400 [2]' =========== -h =-50 h=50 Now sub h=50 into either given equation to find a; i choose [1]; 3(50)+2a=450 2a=300 a=150 So, the cost of the hotel is $50 per day, and the cost of the amusement park is $150 per day.